# Solving hidden monotone variational inequalities with surrogate losses

### v2

@@ 1 Introduction
where the mapping $g:\mathbb{R}^{d}\to\mathcal{Z}\subseteq\mathbb{R}^{n}$ maps model parameters to model outputs, and the set $\mathcal{Z}$ encodes the closure of the set of realizable outputs from the chosen model. Since $F$ is defined over model outputs, it will often be structured, e.g., monotone and smooth.
@@ 2 Background and Related Work
where $f$ is convex-concave. Similar to (2) we can rewrite the problem in parameter space as a constrained problem with respect to model predictions but instead within the VI (1); where $\theta=(\theta^{1},\theta^{2})$, and $g(\theta)=(h^{1}(\theta^{1}),h^{2}(\theta^{2}))$, with operator $F(z)=F(g(\theta))=[\nabla_{z^{1}}f(z^{1},z^{2}),-\nabla_{z^{2}}f(z^{1},z^{2})]^{\top}$ where $z^{1}=h^{1}(\theta^{1})$ and $z^{2}=h^{2}(\theta^{2})$. The solution with respect to model outputs is then equivalent to solving the constrained VI (1). If $F$ is well-conditioned (e.g. smooth and strongly monotone) and $\mathcal{Z}$ is closed and convex, then the projected gradient method $z_{t+1}=\Pi(z_{t}-\eta F(z_{t}))$ converges to a solution $z_{\ast}$ with an appropriate stepsize $\eta$ (Facchinei & Pang, 2003).
@@
Related work.
The surrogate losses proposed by Johnson & Zhang (2020) and Vaswani et al. (2021), apply to supervised learning and RL respectively. [^2] We note that Johnson & Zhang (2020) and Vaswani et al. (2021) provide a more general loss using Bregman divergences. They did not study the VI case, nor do they exploit any convexity properties in the scalar minimization case. Lavington et al. (2023) also study the scalar minimization case and provide convergence to a neighbourhood of a global minimum of (2) and allow for stochasticity. The neighbourhood of convergence depends on an upperbound on the errors $\epsilon_{t}=\|z_{t+1}-z_{t}^{\ast}\|$, where $z_{t}^{\ast}$ is an exact projected gradient step. Therefore, the neighbourhood of convergence scales with the worst error $\epsilon_{t}$ across the trajectory $\{z_{t}=g(\theta_{t})\}_{t\in\mathbb{N}}$. Shrinking the neighbourhood necessitates a double loop algorithm that might spend too much time optimizing the surrogate. Similar results can also be found within the analysis of quasi-Fejér monotone sequences (Franci & Grammatico, 2022). For example, if $z_{t}^{\ast}=T(z_{t})$ where $T$ is a contraction, then $\epsilon_{t}\to 0$ is sufficient to guarantee convergence (see Proposition A.5).
@@
for a proof see Lemma A.2. As expected, the errors are controlled by $\alpha$ and the stepsize $\eta$ in the surrogate. However, the error also scales with the distance to the solution if $F$ is smooth.
@@ 3 Convergence analysis under $\alpha$-descent on surrogate losses
In the VI (1), $\mathcal{Z}$ is convex. There exists a solution within the relative interior, $z_{\ast}\in\operatorname{ri}\mathcal{Z}$. $F$ is both $L$-smooth $\|F(x)-F(y)\|\leq L\|x-y\|$ and $\mu$-strongly monotone $\langle F(x)-F(y),x-y\rangle\geq\mu\|x-y\|^{2}$ for any $x,y\in\mathcal{Z}$ and some $\mu>0$.
@@ 3.1 Convergence and divergence in the deterministic case
Let Assumption 3.1 hold. Then, there exists a fixed stepsize $\eta$ for the the surrogate losses $\{\ell_{t}\}_{t\in\mathbb{N}}$ such that the iterates $\{z_{t}=g(\theta_{t})\}_{t\in\mathbb{N}}$ produced by Algorithm 1 converge linearly to the solution $z_{\ast}$ if either of the following conditions hold:
$1)\;\;\alpha<\frac{\mu}{L}\,,\qquad\text{or}\qquad 2)\;\;\alpha\leq C\eta^{p}\,,\,C,p>0$
@@ 3.2 Unconstrained Stochastic Case
Assumption 3.4 (expected co-coercivity [^4] Note that this is not exactly expected co-coercivity as defined in Loizou et al. (2021) but a direct consequence (Loizou et al., 2021, Lemma 3.4).).
@@
In practice, minimizing ~$\tilde{\ell}_{t}$ exactly is impractical if $n$ is very large due to the sum of squares $\|g(\theta)-g(\theta_{t})\|^{2}$ in (8). Therefore we assume it can be minimized efficiently on expectation with an $\alpha$-expected descent condition (Definition 3.5).
@@
The trajectory $\{\theta_{t}\}_{t\in\mathbb{N}}$ is said to satisfy the $\alpha$-expected descent condition if at each step $t$ the following holds: given the estimate $F_{\xi_{t}}(z_{t})$, the parameter $\theta_{t+1}$ is generated such that
@@
4 A Nonlinear Least Squares Perspective on Surrogate Minimization
@@
The GN method is defined by the update rule $\theta_{t+1}=\theta_{t}-(Dr(\theta_{t})^{\top}Dr(\theta_{t}))^{\dagger}Dr(\theta_{t})^{\top}r(\theta_{t})$. GN inherits the same local quadratic convergence properties as Newton’s method when the Hessian at the minimum $\nabla^{2}f(\theta_{\ast})\approx Dr(\theta_{\ast})^{\top}Dr(\theta_{\ast})$. However, GN is known to struggle with highly non-linear problems, those with large residuals, or if $Dr(\theta_{t})$ is nearly rank-deficient (Björck, 1996). Fortunately, the GN direction is a descent direction of $f$, the DGN method takes steps in the GN direction with a stepsize parameter $\eta_{GN}$ and converges for a sufficiently small stepsize or with line search (Björck, 1996). In cases where $Dr(\theta_{t})$ is nearly rank deficient the LM method can be used.
@@ 4.1 Favourable conditions for gradient descent
Several conditions allow for fast linear convergence of gradient descent for $\ell_{t}(\theta)-\ell_{t}^{\ast}$; see for example  Guille-Escuret et al. (2021). Under linear convergence, $s$-steps of gradient descent guarantees $\alpha=\rho^{s}$ for some $\rho\in[0,1)$, therefore any target value of $\alpha$ is achievable in a finite number of steps that depends only logarithmically in $\alpha$. In this section we show how assumptions used in Sakos et al. (2024) imply two common assumptions for fast convergence of GD: the Polyak-Łojasiewicz (PL) condition (Polyak, 1964; Łojasiewicz, 1963) and $L$-smoothness.
The four assumptions made by Sakos et al. (2024) to study the behaviour of PHGD are: (1-2) $Dg^{\top}$ has both uniformly lower bounded and upper bounded singular values, (3) each component function $g_{i}$ is $\beta$-smooth, (4) and that $\mathcal{Z}$ is bounded. [^5] Although $\mathcal{Z}$ being bounded is not explicitly stated as an assumption it is used within the their main Lemma. In Proposition 4.1 we show that globally lower bounding the singular values of $Dg^{\top}$ guarantees that the composition of a PL function $f$ with $g$ is still PL.
@@
This implies that the surrogate loss $\ell_{t}$ is PL since $\ell_{t}=f_{t}\circ g$ where $f_{t}(z)=\nicefrac{{1}}{{2}}\|z-v_{t}\|^{2}$ is $1$-PL. Similarly, $\ell_{t}$ can be shown to be $L$-smooth if $Dg^{\top}$ has uniformly upper bounded singular values, each $g_{i}$ is $\beta$-smooth, and $\mathcal{Z}$ is bounded (see Lemma B.1). However, these 4 assumptions cannot hold all at once. In Prop. 4.2 we show that if the first three assumptions (1-3) hold, then $\mathcal{Z}$ is unbounded.
@@
This contradiction suggests that the lower bounded singular value assumption is strong. Indeed, it forces the dimension of the parameter space $d$ to be larger than the dimension of the prediction space $n$, and if $g$ is linear it must be surjective $\mathcal{Z}=\mathbb{R}^{n}$. In the case $n=d$ it enforces $Dg$ to be invertible everywhere, which is violated in important cases such as softmax.
@@ 5.2 Minimizing Projected Bellman error
where $v_{\theta}$ is the predicted value vector given parameters $\theta$. Notice that $BE(\theta,\theta^{\prime})$ is the surrogate loss $\nicefrac{{1}}{{2}}\|v_{\theta}-T_{\pi}(v_{\theta^{\prime}})\|^{2}_{\Xi}$. See Appendix D for more details and background.
@@ 5.2.1 Linear Approximation
Figure 3: The average approximation error between GD on the surrogate (13, Surr-GD) and the update (12) are plotted over 10,000 runs in a slow mixing 100-state Markov chain from Bertsekas (2009) and Yu & Bertsekas (2009, Example 3). Surr-GD is observed to converge to the exact update (12) with faster convergence observed as the number of inner steps increase.
@@
The deterministic update is expensive, requiring the full feature matrix $\Phi$. Hence, Bertsekas (2009) constructs an online sequence of estimators $\{\hat{D}_{t}\}_{t\in\mathbb{N}}$, $\{\hat{C}_{t}\}_{t\in\mathbb{N}}$, and $\{\hat{r}_{t}\}$ from a trajectory and suggests the stochastic update (12).
@@
where $\hat{\Xi}$ and $\hat{T}_{\pi}$ correspond to the natural estimators using the empirical distribution (see Appendix  D.1 for more details). Instead of minimizing the loss exactly it can be approximated with GD, thus avoiding the matrix inversion $\hat{D}^{-1}_{t}$, which can make (12) intractable. Figure 3 demonstrates the effectiveness of approximating Bertsekas’ method (12) via GD on the surrogate ~$\tilde{\ell}_{t}$ (13) in a slow mixing 100-state Markov chain.
@@ 5.2.2 Non-linear Projected Bellman Error
For the non-linear setting, we approximate the value function of a fixed reference policy with a two-layer neural network in two Mujoco environments, Ant and Half Cheetah (Todorov et al., 2012). These environments are characterized by continuous state and action spaces, where the agent must control high-dimensional, physics-based simulations to achieve locomotion tasks. The reference policy is derived by training a policy network with the soft-actor-critic algorithm (Haarnoja et al., 2018). For each environment, we consider two different surrogate-based methods with TD(0) as a special case. Similar to the linear case, we aim to find the fixed point (10), and consider the surrogate $\ell_{t}(\theta)=\nicefrac{{1}}{{2}}\|v_{\theta}-T_{\pi}(v_{\theta_{t}})\|_{\Xi}^{2}$. However, due to the large state spaces we must instead consider miniziming the stochastic surrogate ~$\tilde{\ell}_{t}$ (8) from Section 3.2. Notice that ~$\tilde{\ell}_{t}$ consists of two main parts: an inner product term that is linear in $v_{\theta}=g(\theta)$, and a squared error term $\|v_{\theta}-v_{\theta_{t}}\|^{2}$. Our two surrogate methods combine AdamW (Loshchilov, 2017) with two different approaches to approximate these terms.
The first method, Surr-GD, uses the same batch of data to approximate both the inner product part and the squared error term. This fixed batch approach was proposed in Lavington et al. (2023) for scalar minimization, and introduces bias due to approximating the squared error term across all states with a subset of states seen in the batch. Given this approximation we perform either a fixed number of steps (denoted as inner = ##\## steps) or an adaptive number of steps depending on a loss ratio (Thresholded Surr-GD). Taking one gradient step recovers a batch version of TD(0), for more details see Algorithms 2, and 4.
@@
All methods are evaluated with value prediction error, the squared error across a test sample of states visited by a fixed reference policy. The values in these test states are approximated by 100 Monte-Carlo roll-outs. As shown in Figure 4, surrogate methods significantly outperformed TD(0) in both data efficiency (outer loop iterations) and wall-clock time. Interestingly, the surrogate methods converge to a lower prediction error than TD(0), as indicated by the red dashed line being lower than TD(0)’s error curve.

### v3

@@ 1 Introduction
where the mapping $g:\mathbb{R}^{d}\to\mathcal{Z}\subseteq\mathbb{R}^{n}$ maps model parameters to model outputs, and the set $\mathcal{Z}$ encodes the closure of the set of realizable outputs from the chosen model. Since $F$ is defined over model outputs, it will often be structured, e.g., monotone and Lipschitz.
@@ 2 Background and Related Work
where $f$ is convex-concave. Similar to (2) we can rewrite the problem in parameter space as a constrained problem with respect to model predictions but instead within the VI (1); where $\theta=(\theta^{1},\theta^{2})$, and $g(\theta)=(h^{1}(\theta^{1}),h^{2}(\theta^{2}))$, with operator $F(z)=F(g(\theta))=[\nabla_{z^{1}}f(z^{1},z^{2}),-\nabla_{z^{2}}f(z^{1},z^{2})]^{\top}$ where $z^{1}=h^{1}(\theta^{1})$ and $z^{2}=h^{2}(\theta^{2})$. The solution with respect to model outputs is then equivalent to solving the constrained VI (1). If $F$ is well-conditioned (e.g. Lipschitz and strongly monotone) and $\mathcal{Z}$ is closed and convex, then the projected gradient method $z_{t+1}=\Pi(z_{t}-\eta F(z_{t}))$ converges to a solution $z_{\ast}$ with an appropriate stepsize $\eta$ (Facchinei & Pang, 2003).
@@
Surrogate losses in scalar minimization.
The surrogate losses proposed by Johnson & Zhang (2020) and Vaswani et al. (2021), apply to supervised learning and RL respectively. [^2] Note that Johnson & Zhang (2020) and Vaswani et al. (2021) more generally use Bregman divergences. They did not study the VI case, nor do they exploit any convexity properties in the scalar minimization case. Lavington et al. (2023) also study the scalar minimization case and provide convergence to a neighbourhood of a global minimum of (2) and allow for stochasticity. The neighbourhood of convergence depends on an upperbound on the errors $\epsilon_{t}=\|z_{t+1}-z_{t}^{\ast}\|$, where $z_{t}^{\ast}$ is an exact projected gradient step. Therefore, the neighbourhood of convergence scales with the worst error $\epsilon_{t}$ across the trajectory $\{z_{t}=g(\theta_{t})\}_{t\in\mathbb{N}}$. Shrinking the neighbourhood necessitates a double loop algorithm that might spend too much time optimizing the surrogate. Similar results can also be found within the analysis of quasi-Fejér monotone sequences (Franci & Grammatico, 2022). For example, if $z_{t}^{\ast}=T(z_{t})$ where $T$ is a contraction, then $\epsilon_{t}\to 0$ is sufficient to guarantee convergence (see Proposition A.5).
@@
for a proof see Lemma A.2. As expected, the errors are controlled by $\alpha$ and the stepsize $\eta$ in the surrogate. However, the error also scales with the distance to the solution if $F$ is Lipschitz.
@@ 3 Convergence analysis under $\alpha$-descent on surrogate losses
In the VI (1), $\mathcal{Z}$ is convex. There exists a solution within the relative interior, $z_{\ast}\in\operatorname{ri}\mathcal{Z}$. $F$ is both $L$-Lipschitz $\|F(x)-F(y)\|\leq L\|x-y\|$ and $\mu$-strongly monotone $\langle F(x)-F(y),x-y\rangle\geq\mu\|x-y\|^{2}$ for any $x,y\in\mathcal{Z}$ and some $\mu>0$.
@@ 3.1 Convergence and divergence in the deterministic case
Let Assumption 3.1 hold and let $\{z_{t}=g(\theta_{t})\}_{t\in\mathbb{N}}$ be the iterates produced by Algorithm 1. If $\alpha$ and $\eta$ are picked such that $\rho:=1-2\eta(\mu-\alpha L)+(1+\alpha^{2})\eta^{2}L^{2}<1$ then, $z_{t}$ converge linearly to the solution $z_{\ast}$ at the following linear rate:
$\|z_{t+1}-z_{*}\|^{2}\leq\rho^{t}\|z_{1}-z_{*}\|^{2}.$ (8)
Particularly, if $\alpha<\frac{\mu}{L}$ and $\eta<\frac{2(\mu-\alpha L)}{(1+\alpha^{2})L^{2}}$ then $\rho<1$ and if $\alpha\leq\frac{\mu}{2L}$ and $\eta=\frac{2\mu}{5L^{2}}$ then $\rho\leq 1-\frac{\mu^{2}}{5L^{2}}$.
@@
There exists an $L$-Lipschitz and $\mu$-strongly monotone $F$, and a sequence of iterates $\{z_{t}\}_{t\in\mathbb{N}}$ verifying the alpha descent condition with $\alpha<1$ such that $z_{t}$ diverges for any $\eta$.
@@ 3.2 Unconstrained Stochastic Case
Assumption 3.4 (expected co-coercivity [^4] Note that this is not exactly expected co-coercivity but implied by it (Loizou et al., 2021, Lemma 3.4).).
@@
In practice, minimizing ~$\tilde{\ell}_{t}$ exactly is impractical if $n$ is very large due to the sum of squares $\|g(\theta)-g(\theta_{t})\|^{2}$ in (9). Therefore we assume it can be minimized on expectation with the following assumption.
@@
The trajectory $\{\theta_{t}\}_{t\in\mathbb{N}}$ satisfies the $\alpha$-expected descent condition if at for each $t$ the following holds: given $F_{\xi_{t}}(z_{t})$, the parameter $\theta_{t+1}$ is generated such that
@@
Let $\mathcal{Z}=\mathbb{R}^{n}$, and Assumption (3.4) hold. If $F_{\xi}(x)$ is $L$-Lipschitz and $\eta\leq\frac{1}{2(1+c)L}$ where $c\geq 2(1+\alpha^{2})$ then any trajectory $\{\theta_{t}\}_{t\in\mathbb{N}}$ satisfying the $\alpha$-expected descent condition guarantees:
@@
4 A Nonlinear Least Squares Perspective
@@
The GN method is defined by the update rule $\theta_{t+1}=\theta_{t}-(Dr(\theta_{t})^{\top}Dr(\theta_{t}))^{\dagger}Dr(\theta_{t})^{\top}r(\theta_{t})$. GN inherits the same local quadratic convergence properties as Newton’s method when the Hessian at the minimum $\nabla^{2}f(\theta_{\ast})\approx Dr(\theta_{\ast})^{\top}Dr(\theta_{\ast})$. However, GN is known to struggle with highly non-linear problems, those with large residuals, or if $Dr(\theta_{t})$ is nearly rank-deficient (Björck, 1996). Fortunately, the GN direction is a descent direction of $f$, the DGN method takes steps in the GN direction with a stepsize parameter $\eta_{GN}$ and converges for a sufficiently small stepsize or with line search (Björck, 1996). In cases where $Dr(\theta_{t})$ is nearly rank deficient, the LM method can be used.
@@ 4.1 Favourable conditions for gradient descent
Several conditions allow for fast linear convergence of gradient descent for $\ell_{t}(\theta)-\ell_{t}^{\ast}$; see for example  Guille-Escuret et al. (2021). Under linear convergence, $s$-steps of gradient descent guarantees $\alpha=\rho^{s}$ for some $\rho\in[0,1)$, therefore any target value of $\alpha$ is achievable in a finite number of steps that depends only logarithmically in $\alpha$. In this section, we show how assumptions used in Sakos et al. (2024) imply two common assumptions for fast convergence of GD: the Polyak-Łojasiewicz (PL) condition  (Polyak, 1964; Łojasiewicz, 1963, Definition B.2) and $L$-smoothness.
The four assumptions made by Sakos et al. (2024) to study the behaviour of PHGD are: (1-2) $Dg^{\top}$ has both uniformly lower bounded and upper bounded singular values, (3) each component function $g_{i}$ is $\beta$-smooth, (4) and that $\mathcal{Z}$ is bounded.555 The set $\mathcal{Z}$ (note that $\mathcal{Z}$ is the “set of latent variables” denoted as $\mathcal{X}$ in Sakos et al. (2024)) is assumed to be bounded in their Lemma 4 “template inequality”. More precisely, the existence of a constant $D=diam(\mathcal{X})$ (i.e. $\mathcal{X}$ is bounded) is used in equation B.20 in Appendix B. Lemma 4 is then used in Theorems 1-4. In Proposition 4.1 we show that globally lower bounding the singular values of $Dg^{\top}$ guarantees that the composition of a PL function $f$ with $g$ is still PL.
@@
This implies that the surrogate loss $\ell_{t}$ is PL since $\ell_{t}=f_{t}\circ g$ where $f_{t}(z)=\nicefrac{{1}}{{2}}\|z-v_{t}\|^{2}$ is $1$-PL. Similarly, $\ell_{t}$ can be shown to be $L$-smooth if $Dg^{\top}$ has uniformly upper bounded singular values, each $g_{i}$ is $\beta$-smooth, and $\mathcal{Z}$ is bounded (see Lemma B.1). However, these 4 assumptions cannot hold all at once. In Prop. 4.2, we show that if the first three assumptions (1-3) hold, then $\mathcal{Z}$ is unbounded.
@@
This contradiction suggests that the lower bounded singular value assumption is strong. Indeed, it forces the dimension of the parameter space $d$ to be larger than the dimension of the prediction space $n$, and if $g$ is linear then it must be surjective $\mathcal{Z}=\mathbb{R}^{n}$. In the case $n=d$, it enforces $Dg$ to be invertible everywhere, which is violated in important cases such as softmax.
@@ 5.2 Minimizing Projected Bellman error
where $v_{\theta}$ is the predicted value vector given parameters $\theta$. See Appendix D for more details.
@@ 5.2.1 Linear Approximation
Figure 3: The average approximation error between GD on the surrogate (14, Surr-GD) and update (13) over 10,000 runs in a slow mixing 100-state Markov chain from Bertsekas (2009) and Yu & Bertsekas (2009, Example 3). Surr-GD is observed to converge to the exact update (13) with faster convergence for more inner steps.
@@
The deterministic update is expensive, requiring the full feature matrix $\Phi$. Hence, Bertsekas (2009) constructs an online sequence of estimators $\{\hat{D}_{t}\}_{t\in\mathbb{N}}$, $\{\hat{C}_{t}\}_{t\in\mathbb{N}}$, and $\{\hat{r}_{t}\}$ from a trajectory and suggests the stochastic update (13). Although this method applies more generally, in the context of solving (11) update (13) is equivalent to least-squares policy evaluation (Bertsekas & Ioffe, 1996).
@@
where $\hat{\Xi}$ and $\hat{T}_{\pi}$ correspond to the natural estimators using the empirical distribution (see Appendix  D.1 for more details). Instead of minimizing the loss exactly it can be approximated with GD, thus avoiding the matrix inversion $\hat{D}^{-1}_{t}$, which can make (13) intractable. Figure 3 demonstrates the effectiveness of approximating update (13) via GD on the surrogate ~$\tilde{\ell}_{t}$ (14).
@@ 5.2.2 Non-linear Projected Bellman Error
For the non-linear setting, we approximate the value function of a fixed reference policy with a two-layer neural network in two Mujoco environments, Ant and Half Cheetah (Todorov et al., 2012). These environments are characterized by continuous state and action spaces, where the agent must control high-dimensional, physics-based simulations to achieve locomotion tasks. The reference policy is derived by training a policy network with the soft-actor-critic algorithm (Haarnoja et al., 2018). For each environment, we consider two different surrogate-based methods with TD(0) as a special case. Similar to the linear case, we aim to find the fixed point (11), and consider the surrogate $\ell_{t}(\theta)=\nicefrac{{1}}{{2}}\|v_{\theta}-T_{\pi}(v_{\theta_{t}})\|_{\Xi}^{2}$. However, due to the large state spaces we must instead consider miniziming the stochastic surrogate ~$\tilde{\ell}_{t}$ (9) from Section 3.2. Notice that ~$\tilde{\ell}_{t}$ consists of two main parts: an inner product term that is linear in $v_{\theta}=g(\theta)$, and a squared error term $\|v_{\theta}-v_{\theta_{t}}\|^{2}$. Our two surrogate methods use different estimates for the linear and error term.
The first method, Surr-GD, uses the same batch of data to approximate both the inner product part and the squared error term. This fixed batch approach was proposed in Lavington et al. (2023) for scalar minimization, and introduces bias due to approximating the squared error term across all states with states seen in the batch. With this approximation we perform either a fixed number of steps (inner = ##\## steps) or an adaptive number of steps depending on a loss ratio (Thresholded Surr-GD). Taking one step recovers a batch version of TD(0), for more details see Algorithms 2, 4.
@@
All methods are evaluated with value prediction error, the squared error for a test sample of 10,000 states visited by a fixed reference policy. The values in these test states are approximated by 10 Monte-Carlo roll-outs. As shown in Figure 4, surrogate methods significantly outperformed TD(0) in both data efficiency (outer loop iterations) and wall-clock time. Interestingly, the surrogate methods converge to a lower prediction error than TD(0), as indicated by the red dashed line.